OpenStudy (anonymous):

The general value of X in the equation 2sin2x= sinx is: a.nPie+Pie/2 b.npie/4 c. 2npi+pie/2

hartnn (hartnn):

use sin2x =2sin x cos x something gets cancelled out, what ?

OpenStudy (anonymous):

and then? 4sinx cosx = sinx 4cosx=1 and then?

hartnn (hartnn):

be careful, you will consider sin x =0 too 4 sin x cos x = sin x ----> sin x (4cos x-1) = 0 sin x =0 or 4 cosx-1 =0 got this ?

OpenStudy (anonymous):

2sinxcosx=sinx 2cosx=1 cosx=1/2 npi+-pi/6

OpenStudy (anonymous):

theres 2sin2x ..??? not only sin2x

hartnn (hartnn):

can you solve 4cos x-1 = 0 ?

OpenStudy (anonymous):

yeahh

hartnn (hartnn):

well, but with that, you would not get any choice as your answer....because 1/4 is not the cos value of any standard angle...

hartnn (hartnn):

you sure the Q is correct ?

OpenStudy (anonymous):

yeah it was in a book

OpenStudy (debbieg):

Yeah, that factor of 2 in front is definitely a sticky widget here! @rajatgang , be careful, you can NOT divide through by sin(x) in this equation!! Must set is all = 0 and factor. Never divide by an unknown that might be =0.

OpenStudy (anonymous):

im trying it....

OpenStudy (anonymous):

i guess question is wrong ....

OpenStudy (debbieg):

With those as the answer choices? Hmm. Well, books have been known to have errors in them. I dunno, this doesn't look right, and you can punch it into Wolfram and see that none of those are the complete solution set.

OpenStudy (debbieg):

2sin(2x)= sinx Even if you remove the 2 in front on the LHS, you don't get any of those as the complete solution set.

OpenStudy (anonymous):

Here is a step by step solution: http://www.symbolab.com/solutions/trigonometric_equations?query=2sin2x%3D%20sinx